# Plug-and-play imaging with model uncertainty quantification in radio astronomy

### 12 Dec 2023 v1

@@ IV-D Results for image estimation
Visual results for $\Delta T=8\text{h}$ are shown in Figure 3. We first observe that the CLEAN algorithm yields highly blurred and noisy reconstructions. As was observed in [22], the UNet recovers the high intensity results rather well, but struggles at lower intensities (see the extended purple areas): this translates in a relatively good SNR metric, but lower logSNR. Optimisation methods uSARA and SARA show faithful reconstructions, despite noticeable wavelet artefacts around the central point source which is challenging to recover for all algorithms. While no significant difference between the two methods can be noticed in logarithmic scale, we stress that uSARA struggles to recover the high intensity values correctly, which translates in a lower SNR than its constrained counterpart SARA. BM3D based PnP algorithms (i.e. uPnPBM3D and cPnPBM3D) show horizontal and/or vertical artefacts. PnP algorithms with the proposed denoisers (uAIRI and cAIRI) show better reconstruction qualities than other methods, with no significant difference between each method. However, the unconstrained PnP algorithms (3) do not recover high intensities as well as their constrained counterparts (4). This is in line with the results obatined with uSARA and SARA. Interestingly, visual results of OAID and MRID AIRI algorithms are very similar up to small additional artefacts for MRID.
Figure 4 gives the average SNR and logSNR metrics of reconstructions estimated with various algorithms. On these plots, each point is an average over the 15 inverse problems considered for each observation time, error bars representing the 95% confidence intervals. Both uSARA and SARA optimisation algorithms significantly outperform the baseline CLEAN algorithm and the UNet for the two considered metrics. AIRI algorithms improve by 1 to 2 dB over their pure optimisation counterparts uSARA and SARA. In each case (both PnP and pure optimisation), the constrained approach performs better than the unconstrained one, particularly on the SNR metric. We explain this phenomenon by the good ability of constrained methods to recover high intensity values after less iterations than their unconstrained counterparts.
@@
In section III-C, we showed that the peak intensity estimation of the target image affects the choice of the optimal denoising network $\operatorname{D}_{\sigma}$ and the corresponding scaling factor $\beta$ in the uAIRI and cAIRI algorithms. In our simulated experiments, the true peak value in the sought reconstructed image is known (it is equal to 1) and the peak value of the back-projected images vary from 30 to 68 for different measurements. To prove the feasibility of the proposed adaptive peak estimation scheme, we ran the simulated experiments with $\alpha$ set to the exact maximum intensity of the groundtruth image and the maximum intensity of the back-projected image while keeping all other settings the same. Some visual results are shown in Figure 5. Results show that the proposed adaptive peak estimation scheme yields reconstructions with same quality as if the true peak value were known. We also found that if the estimated peak value differs too much from the real peak value, like the situation in Figure 5(c), reconstructions may suffer from chess-board artefacts in the high intensity region. We attribute this suboptimality to the fact that the denoisers are not applied to images with peak value far below 1, a range for which they have not been trained.
@@ IV-E Robustness to denoiser realisation and model uncertainty
In order to investigate the robustness of the proposed approach and following the discussion from Section  III-D, we propose to train $Q=15$ denoisers per training dataset (i.e. OAID and MRID) in the same experimental conditions, but with different random seeds generating the (pseudo) random processes involved during training. This impacts both the initialisation of the denoisers (each being initialised with a different random state) and the random image patches presented to the networks during training, as well as the Gaussian random noise generated in (6). We next plug each denoiser in algorithms (3) and (4). This yields two sets of solutions $\operatorname{Sol}$ of 15 elements, one per training datasets; see Section III-D for more details.
@@
In general, uAIRI tends to provide better logSNR metrics than cAIRI, while the opposite holds for SNR. Furthermore, reconstructions produced with cAIRIOAID and cAIRIMRID exhibit less variation in SNR compared to those generated with uAIRIOAID and uAIRIMRID. Interestingly, both the SNR of the mean image is slightly higher than the mean SNR, indicating that the mean image offers enhanced fidelity. Yet, computing this image necessitates multiple denoiser realizations and algorithm runs.
@@
Figure 8: Imaging of a wide field of view (image size is 2560$\times$2560) containing the the ESO 137-006 galaxy and other radio sources of interest from data acquired with the MeerKAT telescope. (a) imaging result with the baseline CLEAN algorithm; (b) imaging result with the proposed cAIRIOAID algorithm.
@@
Figure 9: Imaging of a wide field of view (image size is 2560$\times$2560) containing the the ESO 137-006 galaxy. Due to the large size of the field of view (same as in Figure 8), we only display a center-crop of size 640$\times$256 over the ESO 137-006 galaxy. (a) shows the back-projected image; (b) the imaging result with the baseline CLEAN algorithm; (c) with the uSARA and (d) SARA algorithm. The white arrow in (d) points towards a filament that not all algorithms manage to recover. The two bottom rows show, from left to right a sample reconstruction with cAIRIOAID in (e), cAIRIMRID in (i), the mean image obtained over 15 reconstructions ((g) and (k)), and the ratio image std/mean in (h) and (l).
@@ V Application to real astronomical data
In this section, we validate the proposed algorithms by imaging of a wide field of view containing the the ESO 137-006 galaxy and other radio sources of interest, from data acquired with the MeerKAT telescope. This field of view has recently sparked interest in the radio astronomy community for the observation of collimated synchrotron threads [58]. We stress that we rely on the same shelf of denoisers that was used in the experiments on simulated data, meaning that no denoisers had to be trained specifically for the experiments in this section.
@@ V-B Imaging results
We show the full imaging field of view in Figure 8 for both the baseline CLEAN algorithm as well as the cAIRIOAID. More reconstruction results are displayed in Figure 9. The maximum intensities of the reconstructions given by various methods are around $0.05$, which yields a dynamic range of approximately $10^{4}$. Therefore, we set $x_{\text{max}}=0.05$ and $a=10^{4}$ in the $\operatorname{rlog}_{a}(\cdot)$ for visualization. Overall, the imaging quality for all algorithms is in line with what was observed on experiments using simulated data (see Section IV). We notice that both the pure optimisation algorithms (SARA and uSARA) and the PnP algorithms significantly improve over CLEAN. However, SARA shows important wavelet artefacts which we explain by calibration errors in the estimation of $H$, and lack of good estimation of the noise standard deviation for $e$ in (1), on which SARA’s hyperparmeters rely (see [60]). Both cAIRIOAID and cAIRIMRID yield images with less artefacts than SARA and better resolution (see bottom-left zoom). Furthermore, notice that the filament shown with the white arrow in Figure 9 (d) is recovered by both cAIRIOAID and cAIRIMRID but not by CLEAN or uSARA. We however underline that neither are completely artefact free: the cAIRIOAID yields reconstructions with dotted artefacts (see the filaments in Figure 9 (e)), while the cAIRIMRID yields light ringing artefacts in the reconstruction (see the filaments in Figure 9 (i)).
@@
We observe different results for denoisers trained on MRID in Figure 9 (i)-(l). The mean image (j) shows a different background value than on the sample (i), along with stronger localized artefacts. The standard deviation map (k) suggests that different denoiser realisations can yield highly different background values in the reconstruction. The ratio map (l) suggests that these variations dominate the signal at very faint intensity.
This study suggests that models trained on the proposed synthetic astronomical dataset (OAID) are more robust than those trained on the synthetic bio dataset (MRID). In practice, we observed that some of the networks trained on the bio dataset showed unstable results when plugged in the PnP algorithm in the case of real imaging, despite low Jacobian spectral norm of the denoisers. We did not observe such phenomenon on the synthetic dataset, and presume this might be related to the null space of the measurement operator in the higher dimensional setting faced in the current real data experiment.
@@ VI Conclusion
Our experiments suggest that cAIRI perform better than its unconstrained counterpart uAIRI, while each respectively improved on their respective pure optimismation counterparts (SARA and uSARA). Our method improves over the state-of-the-art SARA algorithm both in simulated experimental settings, and in real imaging problems. This refines results of our earlier work, where we had pointed the limitations of uAIRI [22]. We believe that this jump in performance is related to the high-dynamic range of the target image, where some point sources with intensities orders of magnitude higher than background pixels are better resolved with constrained algorithms.
By training a shelf of few networks for a wide range of dynamic range of interest in RI imaging, we circumvent the necessity to retrain denoisers when applying AIRI to new imaging setups, where the measurement operator and noise levels may be different. Our work suggests that AIRI algorithms are robust to strong variations in the nature of the training dataset (MRI images versus optical astronomy images), while the capability to adapt the dynamic range of the dataset to the target dynamic range of observation, resulting in the creation of dynamic range-specific denoisers, is a critical feature of the AIRI framework.
Our experiments confirm PnP algorithms’ strong generalization abilities, adapting easily to real data without fine-tuning. In this study, we demonstrate that AIRI denoisers, despite subtle variations in parameter initialisation, noise realisation and patch selection, consistently yield highly similar results when integrated into PnP algorithms. This ensemble of solutions provides an epistemic uncertainty metric for the PnP algorithm. Our results show minimal standard deviations in signal intensity for most denoiser configurations, except for those trained on MRI data. This highlights the need for further fine-tuning when applying bio-trained networks to real imaging scenarios.
@@ Acknowledgments
The authors thank O. Smirnov and A. Dabbech for helpful discussions and for providing calibrated real data. This work was supported by the Engineering and Physical Sciences Research Council (EPSRC). The research used Cirrus, a UK National Tier-2 HPC Service at EPCC funded by the University of Edinburgh and EPSRC (EP/P020267/1). The MeerKAT telescope is operated by SARAO, which is a facility of DST/NRF.

### 14 Dec 2023 v2

@@ IV-D Results for image estimation
Visual results for $\Delta T=8\text{h}$ are shown in Figure 3. We first observe that the CLEAN algorithm yields highly blurred and noisy reconstructions. As was observed in [22], the UNet recovers the high intensity results rather well, but struggles at lower intensities (see the extended purple areas): this translates in a relatively good SNR metric, but lower logSNR. Optimisation methods uSARA and SARA show faithful reconstructions, despite noticeable wavelet artefacts around the central point source which is challenging to recover for all algorithms. While no significant difference between the two methods can be noticed in logarithmic scale, we stress that uSARA struggles to recover the high intensity values correctly, which translates in a lower SNR than its constrained counterpart, SARA. BM3D based PnP algorithms (i.e. uPnPBM3D and cPnPBM3D) show horizontal and/or vertical artefacts. PnP algorithms with the proposed denoisers (uAIRI and cAIRI) show better reconstruction qualities than other methods, with no significant difference between each method. However, the unconstrained PnP algorithms (3) do not recover high intensities as well as their constrained counterparts (4). This is in line with the results obtained with uSARA and SARA. Interestingly, visual results of OAID and MRID AIRI algorithms are very similar up to small additional artefacts for MRID.
Figure 4 gives the average SNR and logSNR metrics of reconstructions estimated with various algorithms. On these plots, each point is an average over the 15 inverse problems considered for each observation time, error bars representing the 95% confidence intervals. Both uSARA and SARA optimisation algorithms significantly outperform the baseline CLEAN algorithm and the UNet for the two considered metrics. AIRI algorithms improve by 1 to 2 dB over their pure optimisation counterparts uSARA and SARA. In each case (both PnP and pure optimisation), the constrained approach performs better than the unconstrained one, particularly on the SNR metric. We explain this phenomenon by the good ability of constrained methods to recover high intensity values after fewer iterations than their unconstrained counterparts.
@@
In section III-C, we showed that the peak intensity estimation of the target image affects the choice of the optimal denoising network $\operatorname{D}_{\sigma}$ and the corresponding scaling factor $\beta$ in the uAIRI and cAIRI algorithms. In our simulated experiments, the true peak value in the sought reconstructed image is known (it is equal to 1) and the peak value of the back-projected images vary from 30 to 68 for different measurements. To prove the feasibility of the proposed adaptive peak estimation scheme, we ran the simulated experiments with $\alpha$ set to the exact maximum intensity of the groundtruth image and the maximum intensity of the back-projected image while keeping all other settings the same. Some visual results are shown in Figure 5, demonstrating that the proposed adaptive peak estimation scheme yields reconstructions with the same quality as if the true peak value were known. We also found that if the estimated peak value differs too much from the real peak value, like the situation in Figure 5(c), reconstructions may suffer from chess-board artefacts in the high intensity region. We attribute this suboptimality to the fact that the denoisers are applied to images with peak value far below 1, a range for which they have not been trained.
@@ IV-E Robustness to denoiser realisation and model uncertainty
In order to investigate the robustness of the proposed approach and following the discussion from Section  III-D, we trained $Q=15$ denoisers per training dataset (i.e. OAID and MRID) in the same experimental conditions, but with different random seeds generating the (pseudo) random processes involved during training. This impacts both the initialisation of the denoisers (each being initialised with a different random state) and the random image patches presented to the networks during training, as well as the Gaussian random noise generated in (6). We next plug each denoiser in algorithms (3) and (4). This yields two sets of solutions $\operatorname{Sol}$ of 15 elements, one per training datasets; see Section III-D for more details.
@@
In general, uAIRI tends to provide better logSNR metrics than cAIRI, while the opposite holds for SNR. Furthermore, reconstructions produced with cAIRIOAID and cAIRIMRID exhibit less variation in SNR compared to those generated with uAIRIOAID and uAIRIMRID. Interestingly, the SNR of both the mean images is slightly higher than the mean SNR, indicating that the mean image offers enhanced fidelity. Yet, computing this image necessitates multiple denoiser realizations and algorithm runs.
@@
Figure 8: Imaging of a wide field of view (image size is 2560$\times$2560) containing the ESO 137-006 galaxy and other radio sources of interest from data acquired with the MeerKAT telescope. (a) imaging result with the baseline CLEAN algorithm; (b) imaging result with the proposed cAIRIOAID algorithm.
@@
Figure 9: Imaging of a wide field of view (image size is 2560$\times$2560) containing the ESO 137-006 galaxy. Due to the large size of the field of view (same as in Figure 8), we only display a center-crop of size 640$\times$256 over the ESO 137-006 galaxy. (a) shows the back-projected image; (b) the imaging result with the baseline CLEAN algorithm; (c) with the uSARA and (d) SARA algorithm. The white arrow in (d) points towards a filament that not all algorithms manage to recover. The two bottom rows show, from left to right a sample reconstruction with cAIRIOAID in (e), cAIRIMRID in (i), the mean image obtained over 15 reconstructions ((g) and (k)), and the ratio image std/mean in (h) and (l).
@@ V Application to real astronomical data
In this section, we validate the proposed algorithms by imaging of a wide field of view containing the ESO 137-006 galaxy and other radio sources of interest, from data acquired with the MeerKAT telescope. This field of view has recently sparked interest in the radio astronomy community for the observation of collimated synchrotron threads [58]. We stress that we rely on the same shelf of denoisers that was used in the experiments on simulated data, meaning that no denoisers had to be trained specifically for the experiments in this section.
@@ V-B Imaging results
We show the full imaging field of view in Figure 8 for both the baseline CLEAN algorithm as well as the cAIRIOAID. More reconstruction results are displayed in Figure 9. The maximum intensities of the reconstructions given by various methods are around $0.05$, which yields a dynamic range of approximately $10^{4}$. Therefore, we set $x_{\text{max}}=0.05$ and $a=10^{4}$ in the $\operatorname{rlog}_{a}(\cdot)$ for visualization. Overall, the imaging quality for all algorithms is in line with what was observed on experiments using simulated data (see Section IV). We notice that both the pure optimisation algorithms (SARA and uSARA) and the PnP algorithms significantly improve over CLEAN. However, SARA shows important wavelet artefacts which we explain by calibration errors in the estimation of $H$, and lack of good estimation of the noise standard deviation for $e$ in (1), on which SARA’s hyperparmeters rely (see [60]). Both cAIRIOAID and cAIRIMRID yield images with fewer artefacts than SARA and better resolution (see bottom-left zoom). Furthermore, notice that the filament shown with the white arrow in Figure 9 (d) is recovered by both cAIRIOAID and cAIRIMRID but not by CLEAN or uSARA. We however underline that neither are completely artefact free: the cAIRIOAID yields reconstructions with dotted artefacts (see the filaments in Figure 9 (e)), while the cAIRIMRID yields light ringing artefacts in the reconstruction (see the filaments in Figure 9 (i)).
@@
We observe different results for denoisers trained on MRID in Figure 9 (i)-(l). The mean image (j) shows a different background value than on the sample (i), along with stronger localized artefacts. The standard deviation map (k) suggests that different denoiser realisations can yield highly different background values in the reconstruction. The ratio map (l) suggests that these variations dominate the signal at very faint intensities.
This study suggests that models trained on the proposed synthetic astronomical dataset (OAID) are more robust than those trained on the synthetic bio dataset (MRID). In practice, we observed that some of the networks trained on the bio dataset showed unstable results when plugged in the PnP algorithm in the case of real imaging, despite the low Jacobian spectral norm of the denoisers. We did not observe such phenomenon on the synthetic dataset, and presume this might be related to the null space of the measurement operator in the higher dimensional setting faced in the current real data experiment.
@@ VI Conclusion
Our experiments suggest that cAIRI perform better than its unconstrained counterpart uAIRI, while each respectively improved on their respective pure optimisation counterparts (SARA and uSARA). Our method improves over the state-of-the-art SARA algorithm both in simulated experimental settings, and in real imaging problems. This refines the results of our earlier work, where we had pointed out the limitations of uAIRI [22]. We believe that this jump in performance is related to the high-dynamic range of the target image, where some point sources with intensities orders of magnitude higher than background pixels are better resolved with constrained algorithms.
By training a shelf of a few networks for a wide range of dynamic range of interest in RI imaging, we circumvent the necessity to retrain denoisers when applying AIRI to new imaging setups, where the measurement operator and noise levels may be different. Our work suggests that AIRI algorithms are robust to strong variations in the nature of the training dataset (MRI images versus optical astronomy images), while the capability to adapt the dynamic range of the dataset to the target dynamic range of observation, resulting in the creation of dynamic range-specific denoisers, is a critical feature of the AIRI framework.
Our experiments confirm PnP algorithms’ strong generalization abilities, adapting easily to real data without fine-tuning. In this study, we demonstrate that AIRI denoisers, despite subtle variations in parameter initialisation, noise realisation and patch selection, consistently yield very similar results when integrated into PnP algorithms. This ensemble of solutions provides an epistemic uncertainty metric for the PnP algorithm. Our results show minimal standard deviations in signal intensity for most denoiser configurations, except for those trained on MRI data. This highlights the need for further fine-tuning when applying bio-trained networks to real imaging scenarios.
@@ Acknowledgments
The authors thank O. Smirnov and A. Dabbech for helpful discussions and for providing calibrated real data. This work was supported by the Engineering and Physical Sciences Research Council (EPSRC), with work funded under the EPSRC grants EP/T028270/1 and EP/T028351/1, and the STFC grant ST/W000970/1. The research used Cirrus, a UK National Tier-2 HPC Service at EPCC funded by the University of Edinburgh and EPSRC (EP/P020267/1). The MeerKAT telescope is operated by SARAO, which is a facility of DST/NRF.